if the speed of sound in helium at 0 degree celsius be 960m/s then what will be the speed in hydrogen at the same temperature given gamma for h2 is 1.40 and helium is 1.67 and the ratio of their molecular masses is 2:1.
 

Grade:12th Pass

2 Answers

Khimraj
5 years ago
speed
Khimraj
5 years ago
speed of sound = \sqrt{\gamma RT/M}
v1/v2 = \sqrt{\gamma_{1} M_{2}/\gamma_{2}M_{1}}
960/v2 = \sqrt{1.4*2/1.67*1}
v2 = 960/1.3
v2 = 738.46 m/s
Hope it clears.